OpenStudy (anonymous):

help with log(base 6) 6 ^x+13 =16 ?

OpenStudy (mertsj):

\[\log_{6}6^{x+13}=16 \]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Yes.

OpenStudy (mertsj):

Can you put it into its exponential form?

OpenStudy (anonymous):

\[(x+13) \log _{6} 6 =16 \] ?

OpenStudy (mertsj):

No. What is the base?

OpenStudy (anonymous):

6.

OpenStudy (mertsj):

What is the exponent?

OpenStudy (anonymous):

x+13

OpenStudy (mertsj):

No. The log is the exponent. What does the equation say the log is?

OpenStudy (anonymous):

OH 6

OpenStudy (mertsj):

No.

OpenStudy (anonymous):

omg 16

OpenStudy (mertsj):